Find three consecutive even integers such that twice the smallest number is twelve more than the largest number?
Answer by CubeyThePenguin(3113)

consecutive even integers: (x-2), x, (x+2)

2(x-2) = 12 + (x+2)
2x - 4 = x + 14
x = 18

integers: 16, 18, 20